Consider a perfectly competitive industry where each firm has a long-run cost function:

C(q) = q3 – 20q2 +120q

Long-run average costs are minimised at an output level of 10 units. The industry demand curve is given by:

Q = 1000 – 40p

where Q represents industry output and p represents the equilibrium price of industry output.

Required:

In the long-run equilibrium, this industry will sustain how many firms?


1
Expert's answer
2021-11-08T09:10:39-0500

"C(q)=q^3-20q^2+120q"

"MC=3q^2-40q+120"

"Q=1000-40P"

To find P,

"10 units= 100-40P"

"\implies P=24.75"

Equate "P=MC"

"24.75=3q^2-40q+120"

"3q^2-40q+95.25=0"

Solving the above quadratic equation, we get:

"q=10.23"

If each perfectly competitive firm is producing 3 units of output and the market output is 10 units, then there will be "\frac{10}{3}=3" firms in the industry.